OpenStudy (moderator):

How many solutions exist for the following system of equations? 2x – 8y = 8 -x + 4y = 4

OpenStudy (anonymous):

zippo

OpenStudy (saifoo.khan):

\[\Huge None!\]

OpenStudy (moderator):

How many solutions exist for the following system of equations? 2x + 3y = 20 -4x - 6y = -28

OpenStudy (anonymous):

has no solution

OpenStudy (saifoo.khan):

lol, none!

OpenStudy (moderator):

for the first or 2nd^

OpenStudy (anonymous):

point is that in these examples one expression on the left hand side is a multiple of the other. but not so on the right

OpenStudy (moderator):

8x + 4y = -24 6x - 8y = -84 is this one zero also?

OpenStudy (anonymous):

U can always plot them to see.. Do u know how to solve simultaneous equations?

OpenStudy (dumbcow):

No this has a solution. notice 8 and 6 are not multiples of each other

OpenStudy (moderator):

oh

OpenStudy (moderator):

so it would be 1 or 2?

OpenStudy (dumbcow):

?? there is one solution a solution is a point (x,y)

OpenStudy (dumbcow):

there are 3 possibilities Zero solutions One solution Infinite solutions

OpenStudy (moderator):

yeah which one

OpenStudy (anonymous):

If u have 2 (straight) lines, they either cross, don't cross or are the same line. Which is it in this case?

OpenStudy (dumbcow):

have you learned elimination method yet?

OpenStudy (moderator):

no not yet

OpenStudy (saifoo.khan):

Oh, elimation work here! Easy way!

OpenStudy (anonymous):

If u have 2 (straight) lines, they either cross, don't cross or are the same line. there are 3 possibilities Zero solutions One solution Infinite solutions

OpenStudy (moderator):

haha lets hear it then saifoo

OpenStudy (saifoo.khan):

Ok, im saying... u hear..

OpenStudy (moderator):

huh

OpenStudy (saifoo.khan):

lol, u wanted to hear it right?

OpenStudy (moderator):

yeah

OpenStudy (moderator):

yeah

OpenStudy (dumbcow):

you should probably stick to graphing them then. just find the intercepts. plug in 0 for x, find y. plug in 0 for y, find x this way you can graph both lines then determine if they cross or not

OpenStudy (anonymous):

Everything u need to know to answer this question (and others similar) is here in this thread, just think about it.